OpenStudy (anonymous):

Please Help! Fan and Medal!!!! Factor: x2 + 5x + 6 A. (x + 3)(x + 2) B. (x - 3)(x + 2) C. (x + 3)(x - 2) D. (x + 5)(x + 1)

OpenStudy (anonymous):

i think its C, but not sure.

OpenStudy (calculusxy):

You can use FOIL to check out which answer choice might work.

OpenStudy (anonymous):

i don't know how to do this

OpenStudy (calculusxy):

So FOIL stands for: First Outside Inner Last

OpenStudy (anonymous):

the answer is A btw

OpenStudy (calculusxy):

Let's take the first choice as an example: \[(x+ 3) (x+2)\] "First": \[x \times x = x^2\] "outside": \[x \times 2 = 2x\] "inner": \[3 \times x = 3x\] "last": \[3 \times 2 = 6\] now put them all together: \[x^2 + 2x + 3x + 6\]

OpenStudy (calculusxy):

Simplify it and then see if it matches :)

OpenStudy (calculusxy):

@kierstenfouts Do you understand how I used FOIL? It's a really useful method!

OpenStudy (anonymous):

i do

OpenStudy (calculusxy):

Great !